A spherical balloon of volume 4.05 x 103 cm3 contains helium at a pressure of 1.15 x 105 Pa. How many moles of helium are in the balloon if the average kinetic energy of the helium atoms is 3.60 x 10-22 J? mol
Answer #1

pressure is given by

P = (2/3) (N/V) E

1.15*10^5 = (2/3)* ( N / 4.05*10^-3)* 3.6*10^-22

N = 1.94*10^24 molecules

number of moles = N / Av No

number of moles = 1.94*10^24 / (6.02*10^23)

n = 3.22 moles

========